B. Suppose that a competitor's product is seen for sale at $10 each. Later, you notice that this same product is now for sale at $12 each. This price has increased by _____ %. C. A training center offers two classes to employees. In Class A, 60 people are currently attending and attendance is growing an average of 6 people per month. In Class B, 80 people are regularly attending and the class is growing at a rate of 2 people per month. Predict when the number of people in each class will be the same.
Answer #1

B. 20%

We see that the original price was $10 and the new price is $12. The percentage increase in price is given by below formula

Hence the price has increased by 20%

C. 5 months

Let x be the number of months when number of people in each classs will be the same. Class A currently has 60 people and number of people are increasung at the rate of 6 per month. Hence after x months the number of people in Class A would be 60+6x.

Class B currently has 80 people and number of people are increasung at the rate of 2 per month. Hence after x months the number of people in Class A would be 80+2x.

To find the value of x, we need to equate the number of people at the end of x months and solve for x. Hence we have the below

60+6x = 80+2x

6x-2x = 80-60

4x = 20

x = 20/4 = 5

Hence the number of people in each class will be the same at month = 5.